Miller Brothers Hardware paid an annual dividend of $0.95 per share last month. Today, the company announced that future dividends will be increasing by 2.6 percent annually. If you require a 13 percent rate of return, how much are you willing to pay to purchase one share of this stock today?
Answer #1

The answer has been presented in the supporting sheet. All the parts has been solved with detailed explanation and calculation. For detailed answer refer to the supporting sheet.

Answer 3 The correct answer is $ 9.37 per share 5 6 7 Purchase price per share = (dividend *(1+growth%))/(return-growth%) =(.
Answer #2

ANSWER :


Given :


D0 = 0.95 ($)

g = 2.6 % =  0.026

r 13% = 0.13


Now, as per DDM at constant growth :


Current value off the stock

= D0 (1 + g) / (r - g)

= 0.95*(1 + 0.026)/ (0.13 - 0.026)

=  9.37 ($)


Maximum that can be paid per share = 9.37 ($) (ANSWER)



answered by: Tulsiram Garg